OpenStudy (anonymous):

Normal Distribution problem: The probability of a plane crashing on a single trip in 1989 was .00000150. Find the approximate probability that in 120,000,000 flights, there will be fewer than 160 crashes. (Round Z to two decimal places. Round your answer to four decimal places.) I'm confused on how to find the mean and the standard deviation in order to get the z score

OpenStudy (kropot72):

The mean is np = 120,000,000 * 0.00000150=180 The standard deviation is \[\sqrt{np(1-p)}=you\ can\ calculate\]

OpenStudy (kropot72):

\[s.d.=\sqrt{179.99973}\]

OpenStudy (kropot72):

@hbaker1 Do you know how to find the z-score?